4x+2y=1, 2x+3y=.70 solve for x and y
Answer by nerdybill(7384)

4x+2y=1, 2x+3y=.70 solve for x and y
.
4x+2y= 1
2x+3y=.70
.
multiply the second equation by -2:
4x+2y= 1
-4x-6y=-1.4
.
add both equations together:
-4y = -0.4
y = 0.1
.
plug the value of y (above) into equation 1 and solve for x:
4x+2y= 1
4x+2(0.1)= 1
4x+0.2= 1
4x = 0.8
x = 0.2
.
(x,y) = (0.2,0.1)